What is the kinetic energy of the electron in a Bohr hydrogen atom in the second excited state?
Answer #1

KE of the Bohr electron

mev2/2 ,

using columbs law and Newton's second law, the centrifugal force = electrostatic force

mev2/r = ke2/r2

Kinetic energy K = ke2/2r

r - Bohr radius is qunatised.

r_n = \frac{n^2\not{h^2}}{m_eke^2} , where n is quantum number 1,2,3, . . . .

the smallest Bohr oribit r_0 = \frac{\not{h^2}}{m_eke^2} = 0.0529 nm

for the second excited state n= 3, n=1 is the ground stated

r3 = 9*ro = 4.761e-10 m

KE = 9.0e-7 *(1.6e-19)2 / 2* 4.761e-10

       = 2.42 e-35 J